1)the standard kilogram is a platinum -iridium cylinder 39.0mm high and 39mm in diameter.what is the density of the material
2) how many grams of copper are required to make a hollow spherical shell having an inner radius of 5.70cm and an outer radius of 5.75cm ?The density of copper is 8.92g/cm³.
3)suppose your hair grows at 1/32inches perday,find the rate at which it grows in nanometes per seconds.
1
Expert's answer
2020-10-30T13:19:44-0400

1) Let us determine the density of the cylinder. The average density is

"\rho = \dfrac{m}{V}" , and in case of the cylinder the volume is

"V=\pi \dfrac{D^2}{4}h=\pi\cdot \dfrac{(3.9\,\mathrm{cm})^2}{4}\cdot3.9\,\text{cm} = 46.6\,\text{cm}^3."

Therefore, the density is

"\rho =\dfrac{1000\,\text{g}}{46.6\,\text{cm}^3} \approx 21.5\,\text{g/cm}^3."


2) First we determine the volume of the shell. It is equal to the difference between the volumes of bigger and smaller spheres:

"V = V_b-V_s = \dfrac43\pi (R_b^3-R_s^3) = \dfrac43\pi (5.75^3-5.70^3) = 20.59\,\text{cm}^3."

Therefore, the mass will be "m = \rho V = 8.92\,\text{g/cm}^3\cdot20.59\,\text{cm}^3 = 183.66\,\text{g}."


3) "1\/32" inch = "1/32\cdot2.54\,\text{cm} = 7.9\cdot10^{-2}\,\mathrm{cm} = 7.9\cdot10^{-4}\,\mathrm{m} = 7.9\cdot10^5\,\mathrm{nm}" per day.

There are "24\cdot3600 = 86400" seconds in a day, so the rate is

"\dfrac{7.9\cdot10^5}{86400} = 9.2" nm/sec.